A continuous random variable M has a pdf given by f(m)={k(1-m/10)for m<or equal to 10

{0,elsewhere

find the value of the constant k,the mean and the variance of x


1
Expert's answer
2021-07-15T05:32:38-0400
"\displaystyle\int_{-\infin}^{\infin}f(m)dm=\displaystyle\int_{0}^{10}k(1-\dfrac{m}{10})dm"

"=k[m-\dfrac{m^2}{20}]\begin{matrix}
 10 \\
 0
\end{matrix}=5k=1"

"k=\dfrac{1}{5}"

"mean=\mu=E[M]=\displaystyle\int_{-\infin}^{\infin}mf(m)dm"

"=\displaystyle\int_{0}^{10}\dfrac{1}{5}m(1-\dfrac{m}{10})dm=[\dfrac{m^2}{10}-\dfrac{m^3}{150}]\begin{matrix}
 10 \\
 0
\end{matrix}=\dfrac{10}{3}"


"E[M^2]=\displaystyle\int_{-\infin}^{\infin}m^2f(m)dm=\displaystyle\int_{0}^{10}\dfrac{1}{5}m^2(1-\dfrac{m}{10})dm"

"=[\dfrac{m^3}{15}-\dfrac{m^4}{200}]\begin{matrix}
 10 \\
 0
\end{matrix}=\dfrac{50}{3}"


"Var(M)=E[M^2]-(E[M])^2"

"=\dfrac{50}{3}-(\dfrac{10}{3})^2=\dfrac{50}{9}"